Each time Caroline goes shopping she decides whether or not to buy fruit.

The probability that she does buy fruit is 0.4.

Independently, she then decides whether or not to buy a CD, with a probability of 0.2 that she does buy a CD.

Work out the probability that she buys fruit or buys a CD or both.



1
Expert's answer
2021-09-27T14:30:59-0400

Probability that she buys fruit = 0.4

Probability that she Does not buys fruit = 1 - 0.4 = 0.6

Probability that she buys CD = 0.2

Probability that she does not buys CD = 1 - 0.2 = 0.8

Probability that she buys fruit or buys a CD or both. = 1 - Probability that she does not but either

= 1 - (0.6)(0.8)

= 1 - 0.48

= 0.52

Or Probability that she buys both = (0.4)(0.2) = 0.08

Probability that she buys fruit or buys a CD or both. = 0.4 + 0.2 - 0.08

= 0.52

Probability that she buys fruit or buys a CD or both. = 0.52